Nicole Clothing manufactures embroidered jackets. It has reported the following standards for direct material and direct labor for the​ year:

Standard DM quantity per​ unit:            3 square feet

Standard cost per sq.​ foot:                                    ​$3.80/sq. ft

Standard DL quantity per​ unit:                2 hours per unit

Standard DL cost per​ hour:                                 ​$9.50/hr

Actual production was​ 14,000 jackets, with​ 33,600 total square feet used at an actual cost of​ $3.90/s                                               q. ft. Total direct labor hours incurred were​ 25,720 at a total cost of​ $249,484. Compute the direct labor rate variance.

A.

​$21,660 F

B.

​$5,144 F

C.

​$21,660 U

D.

​$5,144 U
Answer #1

Labor rate variance

= (SR-AR) *ALH

= (9.50 - 249,484/25,720) * 25,720

= (9.50 - 9.70) * 25,720

= 5144 U